OpenStudy (anonymous):

Can someone please help me and give me a hint on what exactly to do ??? What is the least degree of a polynomial function whose zeros include 5 and 3i? Explain. Plsss help :(

OpenStudy (anonymous):

if the polynomial has a zero of 5, then one factor is \(x-5\)

OpenStudy (kendricklamar2014):

Why didnt you help me @satellite73 :(

OpenStudy (anonymous):

if it has a zero of \(3i\) then it has another zero of \(-3i\) therefore it has two more factors \((x-3i)(x+3i)\)

OpenStudy (anonymous):

if you multiply that out, you get \[(x-3i)(x+3i)=x^2+9\] so your final job is to multiply \[(x-5)(x^2+9)\]

OpenStudy (anonymous):

@KendrickLamar2014 where is your question?

OpenStudy (kendricklamar2014):

Amelia sees two cubic sculptures. She estimates the edge length of the larger sculpture is between 2.5 and 3 times the edge length of the smaller one. How much greater should she estimate the volume of the larger sculpture to be? Explain

OpenStudy (anonymous):

i did answer this \(2.5^3\) to \(3^3\) @KendrickLamar2014

OpenStudy (kendricklamar2014):

@FibonacciChick666 said ir was wrong?

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

@satellite73 I did this other problem and it said to solve the equation and state the type of root and this was the problem x^2+4x+7=0 and I got 2-square root of 11 but idk what the root is,like what does that mean ??

OpenStudy (ranga):

x^2 + 4x + 7 = 0 x = { -4 +/- sqrt(4^2 - 4(1)(7)) } / (2*1) = { -4 +/- sqrt(16-28) } / 2 = { -4 +/- sqrt(-12) } / 2 = { -4 +/- sqrt(-4*3) } / 2 = { -4 +/- 2i * sqrt(3) } / 2 = -2 +/- i * sqrt(3) The roots are: - 2+isqrt(3) and - 2-isqrt(3) Type of roots: They are complex roots.

OpenStudy (anonymous):

Thank you so much

OpenStudy (ranga):

You are welcome.